# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v1

@@ 4 Weighted scoring rules for assessing forecasts of extremes
where $\tilde{y}:=v_{\tau}(y)=\max(y,\tau)$ and $\tilde{F}_{\tau}$ is the censored CDF:
@@ Appendix A Characterizations and properties of the twCRPS
In the following, to simplify the notation, we make frequent use of the so-called chaining function $v_{\tau}$ defined at $v_{\tau}(x):=\max(\tau,x)$.
Characterization 1: equivalence of the twCRPS and the CRPS of a censored distribution.
@@
Let $F$ be a CDF, $y$ an observation, then:
$\text{twCRPS}_{\tau}(F,y)=\text{CRPS}(\tilde{F}_{\tau},\tilde{y})$ (20)
where $\tilde{y}=v_{\tau}(y)=\max(y,\tau)$ and $\tilde{F}_{\tau}$ is the censored CDF, censored at $\tau$:
$\tilde{F}_{\tau}(x)=\begin{cases}0&\text{ if }x<\tau\\ F(x)&\text{ if }x\geq\tau\end{cases}$ (21)
Proof (Lemma A.1).
The proof follows by from the definitions of $v_{\tau}$ and $\tilde{F}$:
[…]
Characterization 2: expression of the twCRPS through the CRPS and a correction factor.
Lemma A.2.
Let $F$ be a CDF, $y$ an observation, then:
$\text{twCRPS}_{\tau}(F,y)=CRPS\left(F,v_{\tau}(y)\right)-\Delta_{F}$ (23)
where $\Delta_{F}$ is a correction factor:
@@
The proof follows by considering integration limits in the twCRPS definition. If $y\leq\tau$ we have:
If $y>\tau$:
@@
Property 1: the twCRPS is stable under censoring.
@@
$\tilde{F}_{t}(x):=\begin{cases}0&\text{ if }x<t\\ F(x)&\text{ if }x>t\end{cases}$ (28)
then if $t\leq\tau$:
$\text{twCRPS}_{\tau}(F_{cens},y)=\text{twCRPS}_{\tau}(F,y)$ (29)
@@
This follows from Lemma A.1. ∎
Property 2: the lower tail twCRPS can be expressed in terms of the CRPS and the upper tail twCRPS.
@@
The lower tail twCRPS (for values under instead of above a threshold):
@@
can be expressed in terms of the upper tail twCRPS:
$\text{twCRPS}_{\tau}^{\text{lower}}+\text{twCRPS}_{\tau}^{\text{upper}}(F,y)=\text{CRPS}(F,y)$ (31)
@@ Appendix B Closed-form expressions for the twCRPS
We make again use of the chaining function $v_{\tau}(x):=\max(\tau,x)$.
Note that the twCRPS is stable under censoring (see Lemma A.3), meaning that all of the closed-form expressions below also apply for censored versions of the distributions.
B.1 Distributions on the real line
B.1.1 Normal
@@
$\text{twCRPS}_{\tau}\left[F,y\right]=|v_{\tau}(y)-y|-v_{\tau}(y)-2\log F(-v_{\tau}(y))+F(\tau)+\log F(-\tau),$ (32)
$\text{twCRPS}_{\tau}\left[F_{\mu,\sigma},y\right]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F,y^{\prime}\right]$ (33)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma}$ and $y^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F$ and $f$ are the standard normal density and CDF respectively.
Source: This equation was given by Wessel et al. [2024].
B.1.2 Logistic
@@
$\text{twCRPS}_{\tau}\left[F,y\right]=-\tau F^{2}(\tau)+v_{\tau}(y)\left[2F(v_{\tau}(y))-1\right]+2\left[f(v_{\tau}(y))-f(\tau)F(\tau)\right]-\frac{1}{\sqrt{\pi}}\left[1-F(\sqrt{2}\tau)\right]$ (34)
$\text{twCRPS}_{\tau}\left[F_{\mu,\sigma},y\right]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F,y^{\prime}\right]$ (35)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma}$ and $y^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F$ and $f$ are the standard logistic density and CDF respectively.
Source: this follows from the expression for the censored logistic distribution in Jordan et al. [2019].
B.1.3 Laplace
@@
$\text{twCRPS}_{\tau}\left[F,y\right]=\begin{cases}|v_{\tau}(y)|+\exp(-|v_{\tau}(y)|)-\frac{3}{4}-\frac{1}{8}\exp(2\tau)&\text{ if }\tau\leq 0\\ |v_{\tau}(y)|+\exp(-|v_{\tau}(y)|)-\tau+\frac{1}{8}\exp(-2\tau)-\exp(-\tau)&\text{ if }\tau>0\\ \end{cases}$ (36)
$\text{twCRPS}_{\tau}\left[F_{\mu,\sigma},y\right]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F,y^{\prime}\right]$ (37)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma}$ and $y^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F$ and $f$ are the standard Laplace density and CDF respectively.
@@
B.1.4 Student’s t
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F_{\nu},y\right]=&-\tau F_{\nu}(\tau)^{2}+v_{\tau}(y)\cdot\left(2F_{\nu}(v_{\tau}(y))-1\right)+2G_{\nu}(\tau)F_{\nu}(\tau)\\ &-2G_{\nu}(v_{\tau}(y))-\bar{B}_{\nu}\left(1-H_{\nu}(\tau)\right)\end{aligned}$ (38)
$\text{twCRPS}_{\tau}\left[F_{\mu,\sigma,\nu},y\right]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F_{\nu},y^{\prime}\right]$ (39)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma}$ and $y^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F_{\nu}$ is the standard Student’s t CDF and $\bar{B}_{\nu},H_{\nu}(x),G_{\nu}(x)$ are defined as in Jordan et al. [2019].
Source: this follows from the expression for the censored Student’s t distribution in Jordan et al. [2019].
B.2 Distributions with positive support
B.2.1 Gamma
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F_{\alpha,\beta},y\right]=$ $-\tau F_{\alpha,\beta}^{2}(\tau)+v_{\tau}(y)[2F_{\alpha,\beta}(v_{\tau}(y))-1]$ (40)
$+\frac{\alpha}{\beta}[1-F_{\alpha,\beta}^{2}(\tau)+2F_{\alpha,\beta}(\tau)F_{\alpha+1,\beta}(\tau)-2F_{\alpha+1,\beta}(v_{\tau}(y))]-\frac{1-F_{2\alpha,\beta}(2\tau)}{\beta B\left(\frac{1}{2},\alpha\right)}$ (41)
where $F$ is the CDF of a Gamma distribution and $B$ the beta function.
@@ B.2.1 Gamma
B.2.2 GPD
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F_{\xi,\mu,\sigma},y\right]=\sigma\left\{|v_{\tau^{\prime}}(y^{\prime})|-\frac{2(1-F_{\xi}(\tau^{\prime}))}{1-\xi}\left[1-\left(1-F_{\xi}\left(\frac{v_{\tau^{\prime}}(y^{\prime})-\tau^{\prime}}{1+\xi\tau^{\prime}/\sigma}\right)\right)^{1-\xi}\right]+\frac{(1-F_{\xi}(\tau^{\prime}))^{2}}{2-\xi}\right\}$ (42)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma}$ and $y^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F$ is the standard logistic density and CDF respectively.
Source: derived by the authors. It follows from the threshold stability of a GPD and the expression for the CRPS of a GPD with point mass in Jordan et al. [2019].
B.2.3 Exponential
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F_{\lambda},y\right]=|v_{\tau}(y)|-\tau+\frac{1}{\lambda}\left[2F_{\lambda}(\tau)-2F_{\lambda}(v_{\tau}(y))+\frac{1}{2}f_{\lambda}(2\tau)\right]$ (43)
where $F$ and $f$ are the exponential distribution and density functions respectively.
Source: derived by the authors. It also follows as special case of the GPD distribution for $\mu=0,\xi=0$, however we give it here explicitly.
B.2.4 Log-normal
twCRPS expression:
$\begin{split}\text{twCRPS}_{\tau}\left[F_{\mu,\sigma},y\right]=&-\tau F^{2}_{\mu,\sigma}(\tau)+v_{\tau}(y)[2F_{\mu,\sigma}(v_{\tau}(y))-1]\\ &+2\exp\left(\mu+\frac{1}{2}\sigma^{2}\right)\left[1-F_{\mu+\sigma^{2},\sigma}(v_{\tau}(y))-I\left(\frac{\log\tau-\mu}{\sigma}-\sigma\right)\right]\end{split}$ (44)
here $F_{\mu,\sigma}$ is the log-normal CDF and $I$ is the integral function:
@@ B.2.4 Log-normal
with $\varphi$ and $\Phi$ the normal density and CDF. This has the statistical interpretation as the probability $P(X>\alpha,Y<X+\beta)$ where $X$ and $Y$ are independent standard normal random variables. We discuss calculation of this integral function in section B.5.
Source: derived by the authors.
B.2.5 Log-Logistic
twCRPS expression:
$\begin{split}\text{twCRPS}_{\tau}\left[F_{\mu,\sigma},y\right]=&v_{\tau}(y)\left[2F_{\mu,\sigma}(v_{\tau}(y))-1\right]-2e^{\mu}B\left(F_{\mu,\sigma}(v_{\tau}(y));1+\sigma,1-\sigma\right)+e^{\mu}(1-\sigma)B\left(1+\sigma,1-\sigma\right)\\ &-\tau\left[1+\ _{2}F_{1}\left(2,\sigma;\sigma+1;-e^{-\mu/\sigma}\tau^{1/\sigma}\right)-2\ _{2}F_{1}\left(1,\sigma;\sigma+1;-e^{-\mu/\sigma}\tau^{1/\sigma}\right)\right]\end{split}$ (46)
where $F_{\mu,\sigma}$ is the log-logistic CDF, $B(\cdot,\cdot),B(\cdot;\cdot,\cdot)$ are the Beta and incomplete Beta functions respectively and ${}_{2}F_{1}$ is the hypergeometric function. The alternative parametrization of the log-logistic distribution in terms of scale $\alpha$ and shape $\beta$ follows from $\mu=\log(\alpha),\sigma=1/\beta$.
@@ B.3 Distributions with flexible support
B.3.1 Truncated Normal $(a,b)$:
twCRPS expression:
$\begin{split}\text{twCRPS}_{\tau}[F^{a,b},y]=&-s\left[\frac{F(\tau)-F(a)}{F(b)-F(a)}\right]^{2}+z\left[2\frac{F(v_{\tau}(y))-F(a)}{F(b)-F(a)}-1\right]\\ &+\frac{2}{F(b)-F(a)}\left[f(v_{\tau}(y))-f(\tau)\frac{F(\tau)-F(a)}{F(b)-F(a)}\right]-\frac{1}{\sqrt{\pi}}\frac{F(\sqrt{2}b)-F(\sqrt{2}\tau)}{\left[F(b)-F(a)\right]^{2}}\end{split}$ (47)
$\text{twCRPS}_{\tau}[F^{a,b}_{\mu,\sigma},y]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F^{a^{\prime},b^{\prime}},y^{\prime}\right]$ (48)
where $\tau^{\prime}:=\tfrac{\tau-\mu}{\sigma},y^{\prime}:=\tfrac{y-\mu}{\sigma},a^{\prime}:=\tfrac{a-\mu}{\sigma}$ and $b^{\prime}:=\tfrac{y-\mu}{\sigma}$. $F$ and $f$ are the standard normal density and CDF respectively.
The case of left truncation at zero (truncation points: $a=0,b=\infty$) follows from $F(b)=F(b^{\prime})=1$ and $a^{\prime}=-\tfrac{\mu}{\sigma}$.
Source: This equation was given by Wessel et al. [2024].
B.3.2 Truncated Logistic $(a,b)$
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F^{a,b},y\right]=&\frac{1}{[F(b)-F(a)]^{2}}\left\{F(\tau)-F(b)+F^{2}(a)\log\frac{F(v_{\tau}(y))}{F(\tau)}\right.\\ &\left.+[1-F(a)]^{2}\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}+F^{2}(b)\log\frac{F(b)}{F(v_{\tau}(y))}+[1-F(b)]^{2}\log\frac{1-F(v_{\tau}(y))}{1-F(b)}\right\}\end{aligned}$ (49)
$\text{twCRPS}_{\tau}[F^{a,b}_{\mu,\sigma},y]=\sigma\cdot\text{twCRPS}_{\tau^{\prime}}\left[F^{a^{\prime},b^{\prime}},y^{\prime}\right]$ (50)
For only left truncation ($b=\infty$) we have:
$\text{twCRPS}_{\tau}\left[F^{a,\infty},y\right]=\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}-\frac{F(a)}{(1-F(a))^{2}}\log F(\tau)-\frac{1+F(a)}{1-F(a)}\log F(v_{\tau}(y))-\frac{1-F(\tau)}{(1-F(a))^{2}}$ (51)
Source: For only left truncation this equation was given by Allen et al. [2021b] and the case for truncation at $(a,b)$ this was derived by the authors.
B.4 Other
B.4.1 Uniform $(0,1)$
twCRPS expression:
$\text{twCRPS}_{\tau}\left[F,y\right]=|y-v_{\tau}(y)|+v_{\tau}(y)^{2}-v_{\tau}(y)+\frac{1}{3}-\frac{\tau^{3}}{3},$ (52)
for $\tau\in(0,1)$.

### v2

@@ 4 Weighted scoring rules for assessing forecasts of extremes
where $\tilde{y}:=v_{\tau}(y)=\max(y,\tau)$ and $\tilde{F}_{\tau}$ is the CDF censored at $\tau$:
@@ Appendix A Characterizations and properties of the twCRPS
In the following, to simplify the notation, we make frequent use of the chaining function $v_{s}$ in its canonical form defined as $v_{s}(x):=\max(s,x)$ for $x\in\mathbb{R}$ and $s\in\mathbb{R}$.
Characterization 1: equivalence of the twCRPS and the CRPS of the censored distribution (lemma 4.1 in section 4).
Lemma 4.1.
Let $F$ be a CDF and $y$ an observation. Then
$\text{twCRPS}_{\tau}(F,y)=\text{CRPS}(\tilde{F}_{\tau},v_{\tau}(y))$ (20)
where $\tilde{F}_{\tau}$ is the CDF censored at $\tau$:
$\tilde{F}_{\tau}(x)=\begin{cases}0,&x<\tau,\\ F(x),&x\geq\tau.\end{cases}$ (21)
Proof.
The proof follows from the definitions of ${\tilde{F}}_{\tau}$ and $v_{\tau}(y)$:
[…]
Characterization 2: expression of the twCRPS through the CRPS and a correction term.
@@
Let $F$ be a CDF and $y$ an observation. Then
$\text{twCRPS}_{\tau}(F,y)=\text{CRPS}\left(F,v_{\tau}(y)\right)-\Delta_{F}$ (23)
where $\Delta_{F}$ is a correction term given as
@@
The proof follows by considering integration limits in the CRPS and twCRPS definitions and the definition of $v_{\tau}(y)$:
@@
The twCRPS is stable under censoring at a point below the threshold $\tau$:
@@
${\tilde{F}}_{t}(x)=\begin{cases}0,&x<t,\\ F(x),&x\geq t.\end{cases}$ (28)
Then for $t\leq\tau$:
$\text{twCRPS}_{\tau}({\tilde{F}}_{t},y)=\text{twCRPS}_{\tau}(F,y)$ (29)
@@
This follows from Lemma 4.1. ∎
The lower-tail twCRPS can be expressed in terms of the CRPS and the upper-tail twCRPS:
@@
The lower-tail twCRPS (for values below instead of above a threshold),
@@
can be expressed in terms of the upper-tail twCRPS as
$\text{twCRPS}_{\tau}^{\text{l}}(F,y)+\text{twCRPS}_{\tau}^{\text{u}}(F,y)=\text{CRPS}(F,y)$ (31)
where $\text{twCRPS}_{\tau}^{\text{u}}$ is equal to $\text{twCRPS}_{\tau}$ as used throughout the paper.
@@ Appendix B Closed-form expressions for the twCRPS
We here provide a list of closed-form expressions for the twCRPS for a couple of standard distributions often used in the context of statistical post-processing. Rather than aiming at an exhaustive list, we restrict ourselves to generic cases of the most important distributions. All of the formulae have been verified against numerical integration. In the limit of the threshold $\tau$ moving towards the lower end of the distribution, known expressions for the CRPS are recovered.
Note that the twCRPS is stable under censoring (see Lemma A.2), meaning that all of the closed-form expressions below also apply to censored versions of the distributions if censoring at a point below the threshold $\tau$.
We again make use of the chaining function $v_{s}(x)=\max(s,x)$ for $x\in\mathbb{R}$ and $s\in\mathbb{R}$.
B.1 Distributions supported on the real line
B.1.1 Normal distribution
@@
$\text{twCRPS}_{\tau}\left(\Phi,y\right)=-\tau\Phi^{2}(\tau)+v_{\tau}(y)\left[2\Phi(v_{\tau}(y))-1\right]+2\left[\varphi(v_{\tau}(y))-\varphi(\tau)\Phi(\tau)\right]-\frac{1}{\sqrt{\pi}}\left[1-\Phi(\sqrt{2}\tau)\right],$ (32)
$\text{twCRPS}_{\tau}\left(F_{\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(\Phi,\frac{y-\mu}{\sigma}\right),$ (33)
where $\varphi$ and $\Phi$ are the standard normal PDF and CDF, respectively, and $F_{\mu,\sigma}(x)=\Phi\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
Source: This formula is given by Wessel et al. [2024].
B.1.2 Laplace distribution
@@
$\text{twCRPS}_{\tau}\left(F,y\right)=\begin{cases}|v_{\tau}(y)|-\frac{3}{4}+\exp(-|v_{\tau}(y)|)-\frac{1}{8}\exp(2\tau),&\tau\leq 0,\\ v_{\tau}(y)-\tau+\exp(-v_{\tau}(y))+\frac{1}{8}\exp(-2\tau)-\exp(-\tau),&\tau\geq 0,\\ \end{cases}$ (34)
$\text{twCRPS}_{\tau}\left(F_{\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F,\frac{y-\mu}{\sigma}\right),$ (35)
where $F(x)=\frac{1}{2}+\frac{1}{2}{\rm sgn}(x)[1-\exp(-|x|)]$ is the standard Laplace CDF, and $F_{\mu,\sigma}(x)=F\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
Source: derived by the authors.
B.1.3 Logistic distribution
@@
$\text{twCRPS}_{\tau}\left(F,y\right)=F(\tau)-1+\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}-\log F(v_{\tau}(y)),$ (36)
$\text{twCRPS}_{\tau}\left(F_{\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F,\frac{y-\mu}{\sigma}\right),$ (37)
where $F(x)=1/[1+\exp(-x)]$ is the standard logistic CDF, and $F_{\mu,\sigma}(x)=F\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
Source: This formula is given by Allen et al. [2021b].
B.1.4 Student’s $t$ distribution
twCRPS expression:
$\text{twCRPS}_{\tau}\left(F_{\nu},y\right)=&-\tau F_{\nu}^{2}(\tau)+v_{\tau}(y)\left[2F_{\nu}(v_{\tau}(y))-1\right]\\ &+2\left[F_{\nu}(\tau)G_{\nu}(\tau)-G_{\nu}(v_{\tau}(y))\right]-\bar{B}_{\nu}\left[1-H_{\nu}(\tau)\right],\end{aligned}$
$\text{twCRPS}_{\tau}\left(F_{\nu,\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{\nu},\frac{y-\mu}{\sigma}\right),$ (38)
where $F_{\nu}$ is the CDF of the Student’s $t$ distribution with $\nu>0$ degrees of freedom, and $F_{\nu,\mu,\sigma}(x)=F_{\nu}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$. Here, $\bar{B}_{\nu}$ and the functions $G_{\nu}$ and $H_{\nu}$ are defined as in Jordan et al. [2019].
Source: derived by the authors. This follows from the CRPS expression for the censored Student’s $t$ distribution in Jordan et al. [2019].
B.2 Distributions with non-negative support
B.2.1 Exponential distribution
twCRPS expression for $\tau\geq 0$:
$\text{twCRPS}_{\tau}\left(F_{\lambda},y\right)=v_{\tau}(y)-\tau+\frac{2}{\lambda}\left[\exp(-\lambda v_{\tau}(y))-\exp(-\lambda\tau)\right]+\frac{1}{2\lambda}\exp(-2\lambda\tau),$ (39)
where $F_{\lambda}$ is the CDF of the exponential distribution with rate parameter $\lambda>0$.
@@ B.2.1 Exponential distribution
B.2.2 Gamma distribution
twCRPS expression for $\tau\geq 0$:
$\text{twCRPS}_{\tau}\left(F_{\alpha,\beta},y\right)=$ $-\tau F_{\alpha,\beta}^{2}(\tau)+v_{\tau}(y)[2F_{\alpha,\beta}(v_{\tau}(y))-1]$
$+\frac{\alpha}{\beta}[1-F_{\alpha,\beta}^{2}(\tau)+2F_{\alpha,\beta}(\tau)F_{\alpha+1,\beta}(\tau)-2F_{\alpha+1,\beta}(v_{\tau}(y))]-\frac{1-F_{2\alpha,\beta}(2\tau)}{\beta B\left(\frac{1}{2},\alpha\right)},$ (40)
where $F_{\alpha,\beta}$ is the CDF of the gamma distribution with shape parameter $\alpha>0$ and rate parameter $\beta>0$, and $B$ denotes the beta function.
@@ B.2.2 Gamma distribution
B.2.3 Log-logistic distribution
twCRPS expression for $\tau\geq 0$ and $10<\sigma<$ 1:
$\begin{split}\text{twCRPS}_{\tau}\left(F_{\mu,\sigma},y\right)=&-\tau F_{\mu,\sigma}^{2}(\tau)+v_{\tau}(y)\left[2F_{\mu,\sigma}(v_{\tau}(y))-1\right]\\ &+2\exp(\mu)\left[B_{\rm u}\left(F_{\mu,\sigma}(v_{\tau}(y));1+\sigma,1-\sigma\right)-B_{\rm u}\left(F_{\mu,\sigma}(\tau);2+\sigma,1-\sigma\right)\right],\end{split}$ (41)
where $F_{\mu,\sigma}$ is the log-logistic CDF with log-location parameter $\mu\in\mathbb{R}$ and log-scale parameter $\sigma>0$, and $B_{\rm u}$ denotes the upper incomplete beta function. The alternative parameterization of the log-logistic distribution in terms of scale parameter $\alpha>0$ and shape parameter $\beta>0$ follows from $\mu=\log\alpha$ and $\sigma=1/\beta$.
Source: derived by the authors.
B.2.4 Log-normal distribution
twCRPS expression for $\tau>0$:
$\begin{split}\text{twCRPS}_{\tau}\left(F_{\mu,\sigma},y\right)=&-\tau F^{2}_{\mu,\sigma}(\tau)+v_{\tau}(y)[2F_{\mu,\sigma}(v_{\tau}(y))-1]\\ &+2\exp\left(\mu+\frac{1}{2}\sigma^{2}\right)\left[1-F_{\mu+\sigma^{2},\sigma}(v_{\tau}(y))-I\left(\frac{\log\tau-\mu}{\sigma}-\sigma,\sigma\right)\right],\end{split}$ (42)
where $F_{\mu,\sigma}$ is the CDF of the log-normal distribution with log-location parameter $\mu\in\mathbb{R}$ and log-scale parameter $\sigma>0$.
Here, $I$ is the integral function
@@ B.2.4 Log-normal distribution
with $\varphi$ and $\Phi$ denoting the standard normal PDF and CDF, respectively. We discuss the evaluation of this integral function in section B.4.
@@ B.3 Distributions with flexible support
B.3.1 Continuous uniform distribution on $[a,b]$
twCRPS expression for $a\leq\tau\leq b$ and $a\leq y\leq b$:
$\text{twCRPS}_{\tau}\left(F,y\right)=\frac{1}{3}\left(1-\tau^{3}\right)+v_{\tau}^{2}(y)-v_{\tau}(y),\qquad 0\leq\tau\leq 1,\qquad 0\leq y\leq 1,$ (44)
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=(b-a)\,\text{twCRPS}_{(\tau-a)/(b-a)}\left(F,\frac{y-a}{b-a}\right),$ (45)
where $F$ is the CDF of the standard uniform distribution, and $F_{a}^{b}(x)=F\left(\frac{x-a}{b-a}\right)$ with $-\infty<a<b<\infty$.